If cos(2m) = .11, which of the following is (cos m)2?

A. 0.555
B. 1.11
C. 2.22
D. .0745
1
Expert's answer
2013-06-24T08:03:47-0400
We know that cos(2m) = 2*( cos(m) )^2 - 1
Then ( cos(m) )^2 = 0.5*( cos(2m) + 1 ) = 0.5* ( 0.11 + 1 )=0.5*1.11=0.555
so correct answer ( cos(m) )^2 = .555